A 9.1 kg bowling ball traveling at 19.1 m/s strikes a 1.6 kg bowling pin. The impact slows the bowling ball to 6.9 m/s.



Determine the velocity given to the pin.
1
Expert's answer
2020-01-31T10:10:47-0500
"P=M(v_i-v_f)"

"u=\frac{M}{m}(v_i-v_f)"

"u=\frac{9.1}{1.6}(19.1-6.9)=69\frac{m}{s}"